Circular No. 5991

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET SHOEMAKER (1994k) Carolyn S. Shoemaker reports her discovery of another comet on Palo-

mar 0.46-m Schmidt films. The following positions are available:

1994 UT α_{2000} δ_{2000} m_{1} Observer
May 14.41284 17^{h}21^{m}58.^{s}03 − 9^{o}43′26′′.5 17.3 Shoemaker
14.44739 17 21 56.96 − 9 43 34.6  _{”}
20.65317 17 18 41.21 −10 10 30.8 17.5 Kojima
20.66286 17 18 40.86 −10 10 33.9  _{”}
20.68248 17 18 40.22 −10 10 38.6  _{”}

E. M. Shoemaker, C. S. Shoemaker, D. H. Levy, and T. B. Spahr (Palomar).

Films scanned and measured by C. S. Shoemaker. T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/6 reflector + CCD.

Comet is diffuse with strong condensation. Communicated by S. Nakano.

SECOND BATSE BURST CATALOGUE C. Meegan, Marshall Space Flight Center, reports for the BATSE team:

“The second BATSE catalogue of γ-ray bursts is now available through the Gamma-Ray Observatory Science Support Center. The catalogue comprises 585 bursts observed through 1993 Mar. 9 and includes the previously released 1B catalogue. To access the catalogue, INTERNET users should type ‘telnet grossc.gsfc.nasa.gov’; DECNET users should type ‘set host grossc’. The login name is ‘gronews’, and no password is required.”

SUPERNOVA 1994I IN NGC 5194 R. Rudy and Y. Dotan, The Aerospace Corporation; and C. Wood-

ward, J. Cole, and T. Hodge, University of Wyoming, report the following near-infrared magnitudes (with 1σ errors of 0.08 mag) for SN 1994I, obtained on Apr. 23.2 UT using the Aerospace near-infrared camera (which incorporates a NICMOS3 array) on the 2.3-m telescope of the Wyoming Infrared Observatory: J = 13.18; H = 12.70; [2.00–2.32 μm] = 12.33.

SUPERNOVA 1994P IN UGC 6983 Corrigendum. On IAUC 5989, line 1, delete Princeton University; line

5, for aligned read nearly aligned [the major axis of the galaxy is aligned almost exactly east-west]

1994 May 20 Daniel W. E. Green